OpenStudy (anonymous):

If f(x) = −4x − 1, find f(−2). −9 −7 −10 7 Question 3 If h(x) = 2x + 3, find the value of x if h(x) = 9. 21 3 6 9 Question 4 Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. Choose the cost function that represents this scenario if x equals the number of nights Laura has the movie. c(x) = 2.00x + 0.50 c(x) = 2.00 + 0.50x c(x) = 2.50x c(x) = (2.00 + 0.50)x Question 5 If g(x) = 2(x − 4), find the value of x if g(x

OpenStudy (anonymous):

FOR ques no1 f(x)=-4x-1 =>f(-2)=-4.-2-1 =8-1 =7

OpenStudy (nurali):

f(x) = −4x − 1 f(−2)=-4(-2)-1 f(−2)=8-1 f(−2)=7 h(x) = 2x + 3 9=2x+3 2x=9-3 2x=6 x=6/2 x=3

OpenStudy (nurali):

c(x) = 2.00x + 0.50 means 50 cents flat fee and 2.00 per night c(x) = 2.50x means 2.50 per night and 0 flat fee c(x) = (2.00 + 0.50)x means 2.50 per night as well

OpenStudy (nurali):

Question 5 in complete

OpenStudy (nurali):

tell me last question

OpenStudy (nurali):

Reply